# A uniformity principle for spatial matching

### v1

@@ 1 Introduction
To answer this question, we model the system as a bipartite random geometric graph [Pen03]. There are $n$ supply nodes and $m$ demand nodes embedded in a feature space, taken to be the unit hypercube with dimension $k\geq 1$. The features of demand and supply nodes are drawn independently from the uniform distribution. A supply node is compatible to serve a demand node if the demand node lies within a ball centered at the supply node; this ball is the supply node’s service region with volume equal to its service range. This construction induces a compatibility graph that is a bipartite geometric graph, with edges connecting each supply node to the demand nodes that lie within its service region. We assume each supply node can serve at most one demand node. We therefore measure the effectiveness of a service-range allocation by the expected maximum number of demand nodes that can be served, equivalently the expected size of a maximum matching in the induced compatibility graph.
@@ 1.1 Related literature
When service ranges are unconstrained on $0,1]^{k}$, the bipartite matching problem reduces to Euclidean optimal transport (OT) between two i.i.d. uniform samples. Classical results show that the optimal total transportation cost (sum of Euclidean distances) grows with $n$ (e.g., $\Theta(\sqrt{n})$ for $k=1$, $\Theta(\sqrt{n\log n})$ for $k=2$, and $n^{1/(k+1)}$ for $k\geq 3$ [AKT84, Tal94]. In contrast, our model forces locality through the service range, yielding a total “cost” of $\Theta(1)$. Rather than fulfilling all possible demand with unbounded transportation distances, we characterize how to optimally allocate a fixed budget to maximize expected matches, showing that uniform allocation dominates and that matching sizes increase with the total budget. This bounded-range regime captures a practical effect, where platforms let some demand go unfulfilled (or rolled on to the next batch), rather than serving them at a prohibitive cost.
@@ 1.2 Model
$\|s_{i}-d_{j}\|_{2}\leq\left({r_{i}}/{n}\right)^{1/k}.$
Equivalently, supply node $i$ serves only demand nodes within a ball of volume $r_{i}/n$ centered at $s_{i}$, where the scaling by $n$ ensures operation in the sparse regime of random geometric graphs with $\Theta(1)$ degree per supply node on average. Let
@@
In our model, the service range parametrizes the volume of the service region, which equals to the expected number of compatible demand nodes. An alternative choice is to have the service range parameterize the radius of the service region, so that supply node $i$ can serve demand node $j$ if $\|s_{i}-d_{j}\|_{2}\leq r_{i}/n$. This interpretation is natural in applications such as ride-hailing, where distance has a direct operational meaning. We study this alternative specification through simulation in Section 4 and show that, for any $k\geq 2$, it leads to qualitatively different behavior, as the expected number of compatible demand nodes then grows superlinearly in the radius.
@@ 1.3.1 Uniformity principle
$\sum_{i=1}^{k}x_{i}^{\downarrow}\geq\sum_{i=1}^{k}y_{i}^{\downarrow}\quad(k=1,\dots,n-1),\qquad\sum_{i=1}^{n}x_{i}^{\downarrow}=\sum_{i=1}^{n}y_{i}^{\downarrow}.$
@@ 1.4 Notation and paper organization
The rest of the paper is organized as follows. In Section 2, we establish the uniformity principle (Theorem 1); the proof in one dimension is given in Section 2.1, and the extension to higher dimensions is developed in Section 2.2. Section 3 then studies the dual-service range model and proves Theorems 2 and 3. Simulation results are presented in Section 4, and concluding remarks appear in Section 5. Proofs of supporting lemmas and auxiliary results are deferred to the Appendix.
@@ 2 Analysis for the uniformity principle
In this section we prove the uniformity principle (Theorem 1), with technical lemmas stated here whose proofs are deferred to Section B. We analyze $k=1$ and $k\geq 2$ separately because the underlying RGGs exhibit a structural dichotomy: when $k=1$, all connected components are $O(\log n)$ with high probability, while for $k\geq 2$ a giant component of size $\Theta(n)$ emerges once the average degree crosses a constant threshold ([Pen03, Chs. 10–11]). The $k=1$ proof localizes to small components and yields sharper bounds, whereas the argument for $k\geq 2$ controls long-range dependencies induced by the giant component, leading to a slightly loose bound.
@@ 2.1 Analysis for $k=1$
Let $\left\{s_{i}\right\}_{i=1}^{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathbb{D}_{1}$ and $\left\{d_{j}\right\}_{j=1}^{m}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathbb{D}_{2}$ be mutually independent. Assume that $\mathbb{D}_{1}$ and $\mathbb{D}_{2}$ have densities $g_{1}$ and $g_{2}$ respectively on $[0,1]$, that are both $\eta$-Lipschitz and bounded in $1/\eta,\,\eta]$ for some $\eta\geq 1$. For any $\gamma\geq\max\{1,\xi^{-1}\}$ and any two service-range vectors $\mathsf{R},\,\mathsf{R}^{\prime}\in[0,\,\gamma]^{n}$ satisfying
@@ 2.2 Analysis for $k\geq 2$
Only demand nodes whose neighbors are within distance $(\gamma/n)^{1/k}$ of a cell boundary are affected by trimming. Then,
@@
where each constant depends on the variables in the subscript only.
@@ 2.2.1 Proof of Proposition 5
The proof follows the same structure as that of Proposition 2. We use the same notation, generalizing $\delta_{m}(r,\mathsf{R})$ to $\delta^{\varepsilon}_{m}(r,\mathsf{R})$ and $\rho_{m}^{\mathsf{R}}(r_{x},r_{y})$ to $\rho_{m}^{\mathsf{R},\varepsilon}(r_{x},r_{y})$ by replacing $G$ with $G_{\varepsilon}$ in the definitions. The main technical challenge is to show the success probability function $\delta_{m}^{\varepsilon}(\cdot,\mathsf{R})$ is strongly concave.
@@
For supply nodes $x=(s_{x},r_{x})$ and $y=(s_{y},r_{y})$ with $s_{x},s_{y}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Uniform}([0,1]^{k})$ independent of $G\sim\mathbb{G}(m,\mathsf{R})$ with $\mathsf{R}\in{\mathbb{R}}_{\geq 0}^{n-2}$, then
$\rho_{m}^{\mathsf{R},\varepsilon}(r_{x},r_{y})\leq\beta_{k,\gamma}\,\varepsilon^{-k}/{n}\,,\text{ where }\beta_{k,\gamma}\triangleq\left(2k\right)^{k}\gamma$
@@ 4. $\mathcal{M}^{\prime}\leftarrow\emptyset$ $\triangleright$ matching
5. for $k=1,\cdots,|\mathsf{D}|$ do
6.   if $d_{(k)}$ has a neighbor in $\mathsf{S}^{\prime}$ then
7.    $k^{*}\leftarrow\mathop{\arg\min}\{s_{i}+r_{i}/n:s_{i}\text{ is a neighbor of $d_{(k)}$ in $\mathsf{S}^{\prime}$}\}$ $\triangleright$ greedy matching rule
8.    $\mathsf{S}^{\prime}\leftarrow\mathsf{S}^{\prime}-s_{k^{*}}$
9.    $\mathsf{D}^{\prime}\leftarrow\mathsf{D}^{\prime}-d_{(k)}$
10.    $\mathcal{M}^{\prime}\leftarrow\mathcal{M}^{\prime}+(d_{(k)},s_{k^{*}})$
@@ 4 Simulations
In this section, we present simulations illustrating (i) the uniformity principle for service ranges, (ii) the failure of an analogous monotonicity when one works with service radii, and (iii) the agreement of the Markov chain formula (2) versus the true maximum matching.
Throughout, we generate $n=400$ supply nodes and $n=400$ demand nodes i.i.d. uniformly on $0,1]^{k}$, form the induced bipartite graph under the specified connectivity rule, and compute the maximum matching size. Curves report the average matching fraction over $10^{4}$ independent trials, and shaded regions indicate $\pm 1$ standard deviation over $10^{4}$ independent trials.
To generate service range vectors that are ordered according to majorization, we use a randomized version of our dual service-range model, parameterized by a scalar $\alpha\in[0,1]$:
$\mathsf{R}_{i}(\alpha)=r(\alpha)+b(\alpha)\,X_{i}(\alpha),~~~~~~X_{i}(\alpha)\stackrel{{\scriptstyle\mathrm{i.i.d}}}{{\sim}}\mathrm{Bernoulli}\big(p(\alpha)\big)\,$,
i.e. each supply node has service range $r(\alpha)$ with probability $1-p(\alpha)$, and range $r(\alpha)+b(\alpha)$ with probability $p(\alpha)$. The parameters $r(\alpha),\,b(\alpha),\,p(\alpha)$ are chosen so that the mean service range $\bar{r}$ is fixed independent of $\alpha$, i.e.
$\bar{r}\triangleq\mathbb{E}[\mathsf{R}_{i}(\alpha)]=r(\alpha)+p(\alpha)\,b(\alpha)\,.$
For a two-point distribution, increasing the variance while keeping the mean fixed produces a more unequal allocation, i.e. a more majorized vector. Accordingly, we consider three one-parameter families in which the variance increases monotonically with $\alpha$. The resulting parametrizations (fixed $r$, fixed $b$ and fixed $p$) are summarized in Table 2.
Family Fixed Varied in $\alpha$ Implied by $\bar{r}=r+pb$
Fixed $r$ $r(\alpha)=r_{0}$ $p(\alpha)=p_{\max}-\alpha\big(p_{\max}-p_{\min}\big)$ $b(\alpha)=(\bar{r}-r_{0})/p(\alpha)$
Fixed $b$ $b(\alpha)=b_{0}$ $p(\alpha)=p_{\min}+\alpha\big(p_{\max}-p_{\min}\big)$ $r(\alpha)=\bar{r}-p(\alpha)b_{0}$
Fixed $p$ $p(\alpha)=p_{0}$ $r(\alpha)=r_{\max}-\alpha\big(r_{\max}-r_{\min}\big)$ $b(\alpha)=(\bar{r}-r(\alpha))/p_{0}$
@@
We evaluate the uniformity principle by choosing $\mathsf{R}(\alpha)$ so that
$\alpha_{1}<\alpha_{2}\ \implies\ \mathsf{R}(\alpha_{1})\preceq\mathsf{R}(\alpha_{2}),$
and then plotting the matching fraction against $\alpha$. For the three families (fixed $r$, fixed $b$, fixed $p$) defined in Table 2, the parameter choices used to generate Figure 11 are reported in Table 3.
Dimension $k$ Family Mean/fixed parameter(s) Varying range
11 Fixed $r$ $\bar{r}=1,\ r_{0}=0.5$ $p_{\min}=0.2,\ p_{\max}=0.8$
Fixed $b$ $\bar{r}=1,\ b_{0}=0.6$ $p_{\min}=0.01,\ p_{\max}=0.5$
Fixed $p$ $\bar{r}=1,\ p_{0}=0.3$ $r_{\min}=0.5,\ r_{\max}=1$
$2,3$ Fixed $r$ $\bar{r}=0.0012,\ r_{0}=0.0002$ $p_{\min}=0.2,\ p_{\max}=0.8$
Fixed $b$ $\bar{r}=0.0012,\ b_{0}=0.0022$ $p_{\min}=0.0,\ p_{\max}=0.5$
Fixed $p$ $\bar{r}=0.0012,\ p_{0}=0.5$ $r_{\min}=0.0005,\ r_{\max}=0.001$
Table 3: Numerical choices used to generate $\mathsf{R}(\alpha)$ in Figure 11. In each row, the remaining parameter is determined by the mean constraint $\bar{r}=r+pb$, and the dependence on $\alpha$ is linear as specified in Table 2. (The same parameters are used for $k=2$ and $k=3$.)
Figure 11 shows that in all three families and in all tested dimensions, the matching fraction is monotonically decreasing in $\alpha$, consistent with the uniformity principle: smaller $\alpha$ produces a less majorized (more equal) service range vector and leads to better matching performance.
@@
Figure 11: Illustration of the uniformity principle. Each panel shows three curves corresponding to the fixed $r$ (fixed $r$), fixed $b$ (fixed $b$), and fixed $p$ families.
Observations on radius modeling alternative.
As discussed in Remark 1, we now explore an alternative modeling choice in which service range parametrizes the radius of the service region rather than its volume. We focus on the case of dimension $2$, where the distinction between the two specifications is particularly transparent. The two models are instantiated as follows.
  • (Service range parametrizes volume) Given a vector $\mathsf{R}$ of service ranges, $s_{i}$ and $d_{j}$ are connected iff
    $\|s_{i}-d_{j}\|_{2}\leq\sqrt{\mathsf{R}_{i}/(\pi\,n)}.$
  • (Service range parametrizes radius) Given a vector $\mathsf{R}$ of service ranges, $s_{i}$ and $d_{j}$ are connected iff
    $\|s_{i}-d_{j}\|_{2}\leq\mathsf{R}_{i}/n.$
We reuse the same three one-parameter families $\mathsf{R}(\alpha)$ from Table 2 (fixed $r$, fixed $b$, fixed $p$), but implement $\mathsf{R}(\alpha)$ either as a vector of volumes, or as a vector of radii of the service region. Table 4 lists the numerical choices used under each interpretation. Figure 12 reports the resulting matching fractions (averaged over $10^{4}$ trials, with $\pm 1$ standard deviation shading). As established in Theorem 1, the model where service range parametrizes volume obeys the uniformity principle (monotone decrease in $\alpha$); Figure 12 verifies this and also shows that the radius model does not obey an analogous monotonicity. In particular, one can find $\alpha_{1}<\alpha_{2}$ with $\mathsf{R}(\alpha_{1})\preceq\mathsf{R}(\alpha_{2})$ yet a larger matching fraction at $\alpha_{2}$ under the radius interpretation.
(a) Fixed $r$
(b) Fixed $b$
(c) Fixed $p$
Figure 12: Radius vs. range in $k=2$ under the three one-parameter families.
(a) Versus $r$
(b) Versus $b$
(c) Versus $p$
@@
Parametrization Radius interpretation Volume interpretation
Fixed $r$ $\bar{r}=0.015,\,r_{0}=0,\,p\in[0.2,0.8]$ $\bar{r}=0.001,\,r_{0}=0,\,p\in[0.2,0.8]$
Fixed $b$ $\bar{r}=0.02,\,b_{0}=0.04,\,p\in[0,0.5]$ $\bar{r}=0.0012,\,b_{0}=0.0007,\,p\in[0,0.5]$
Fixed $p$ $\bar{r}=0.02,\,p_{0}=0.5,\,r\in[0.001,0.018]$ $\bar{r}=\tfrac{0.02}{12},\,p_{0}=0.5,\,r\in[\tfrac{0.001}{12},\,\tfrac{0.018}{12}]$
@@
Figure 13 checks the accuracy of the closed-form expression (2), which is derived from the stationary distribution of the Markov chain $\boldsymbol{\psi}(t)$. For each parameter setting, we compute (i) the matching fraction predicted by (2) and (ii) the exact matching fraction obtained by solving the underlying bipartite matching instance via the Hopcroft–Karp algorithm. We plot the two quantities under three separate parameter sweeps, varying $r$, $b$, and $p$ (holding the remaining parameters fixed). Across all sweeps, the predicted and exact curves agree, empirically validating (2).
Bounds for the dual-service range model.
In Remark 4, we derived complementary upper and lower bounds on the fraction of matched nodes in the dual-service range model. The upper bound (7) follows from the uniformity principle, while the lower bounds (8) and (9) are obtained by splitting the random graph into two subgraphs and lower bounding the maximum matching size via matchings within each part. Figure 14 compares these bounds across a range of parameter values by plotting them as functions of $r$, $b$, and $p$ (with the remaining parameters held fixed).

### v2

@@ 1 Introduction
To answer this question, we model the system as a bipartite random geometric graph [Pen03]. There are $n$ supply nodes and $m$ demand nodes embedded in a feature space, taken to be the unit hypercube with dimension $k\geq 1$. The features of demand and supply nodes are drawn independently from the uniform distribution. A supply node is compatible to serve a demand node if the demand node lies within a ball centered at the supply node; this ball is the supply node’s service region with volume determined by its service range. This construction induces a compatibility graph that is a bipartite geometric graph, with edges connecting each supply node to the demand nodes that lie within its service region. We assume each supply node can serve at most one demand node. We therefore measure the effectiveness of a service range allocation by the expected maximum number of demand nodes that can be served, equivalently the expected size of a maximum matching in the induced compatibility graph.
@@ 1.1 Related literature
When service ranges are unconstrained on $0,1]^{k}$, the bipartite matching problem reduces to Euclidean optimal transport (OT) between two i.i.d. uniform samples. In particular, if the transportation cost is the sum of Euclidean distances raised to the power $k$, $\min_{\sigma\in S_{n}}\sum_{i=1}^{n}\|X_{i}-Y_{\sigma(i)}\|^{k}\;=\;n\,\mathsf{OT}_{n,k},$ then classical results imply the expected optimal total cost scales as $\mathbb{E}\!\left[n\,\mathsf{OT}_{n,k}\right]=\Theta(\sqrt{n})$ for $k=1$, $\mathbb{E}\!\left[n\,\mathsf{OT}_{n,k}\right]=\Theta(\log n)$ for $k=2$, and $\mathbb{E}\!\left[n\,\mathsf{OT}_{n,k}\right]=\Theta(1)$ for $k\geq 3$ [AKT84, Tal94]. In contrast, our model forces locality through the service range, yielding a total “cost” of $\Theta(1)$ for all $k\geq 1$. Rather than fulfilling all possible demand with unbounded transportation distances, we characterize how to optimally allocate a fixed budget to maximize expected matches, showing that uniform allocation dominates and that matching sizes increase with the total budget. This bounded-range regime captures a practical effect, where platforms let some demand go unfulfilled (or rolled on to the next batch), rather than serving them at a prohibitive cost.
@@ 1.2 Model
$\|s_{i}-d_{j}\|_{2}\,\leq\,\left({r_{i}}/{n}\right)^{1/k}\,$,
Equivalently, supply node $i$ serves only demand nodes within a ball of volume $\kappa_{k}\,r_{i}/n$ centered at $s_{i}$, where $\kappa_{k}=\pi^{k/2}/\Gamma(1+k/2)$ is the volume of a unit ball in $\mathbb{R}^{k}$, and the scaling by $n$ ensures operation in the sparse regime of random geometric graphs with $\Theta(1)$ degree per supply node on average. Let
@@
In our model, the service range parametrizes the volume of the service region, which is proportional to the expected number of compatible demand nodes. An alternative choice is to have the service range parametrize the radius of the service region, so that supply node $i$ can serve demand node $j$ if $\|s_{i}-d_{j}\|_{2}\leq r_{i}/n^{1/k}$, where the scaling by $n^{1/k}$ ensures a $\Theta(1)$ degree per supply node on average. This interpretation is natural in applications such as ride-hailing, where distance has a direct operational meaning. Note that these two parametrizations are equivalent when $k=1$. We study the radius model through simulation in Section 4 and show that for $k=2$, it leads to qualitatively different behavior.
@@ 1.3.1 Uniformity principle
$\sum_{i=1}^{\ell}x_{i}^{\downarrow}\geq\sum_{i=1}^{\ell}y_{i}^{\downarrow}\quad(\ell=1,\dots,n-1),\qquad\sum_{i=1}^{n}x_{i}^{\downarrow}=\sum_{i=1}^{n}y_{i}^{\downarrow}.$
@@ 1.4 Notation and paper organization
The rest of the paper is organized as follows. In Section 2, we establish the uniformity principle (Theorem 1); the proof in one dimension is given in Section 2.1, and the extension to higher dimensions is developed in Section 2.2. Section 3 then studies the dual-service range model and proves Theorems 2 and 3. Simulation results are presented in Section 4, and concluding remarks appear in Section 5. Proofs of supporting lemmas and auxiliary results are deferred to the appendices.
@@ 2 Analysis for the uniformity principle
In this section we prove the uniformity principle (Theorem 1), with technical lemmas stated here whose proofs are deferred to Appendix B. We analyze $k=1$ and $k\geq 2$ separately because the underlying RGGs exhibit a structural dichotomy: when $k=1$, all connected components are of size $O(\log n)$ with high probability, while for $k\geq 2$ a giant component of size $\Theta(n)$ emerges once the average degree crosses a constant threshold ([Pen03, Chs. 10–11]). The $k=1$ proof localizes to small components and yields sharper bounds, whereas the argument for $k\geq 2$ controls long-range dependencies induced by the giant component, leading to a slightly loose bound.
@@ 2.1 Analysis for $k=1$
Let $\left\{s_{i}\right\}_{i=1}^{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathbb{D}_{1}$ and $\left\{d_{j}\right\}_{j=1}^{m}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathbb{D}_{2}$ be mutually independent. Assume that $\mathbb{D}_{1}$ and $\mathbb{D}_{2}$ have densities $g_{1}$ and $g_{2}$ respectively on $[0,1]$, that are both $\eta$-Lipschitz and bounded in $1/\eta,\,\eta]$ for some $\eta\geq 1$. For any $\gamma\geq\max\{1,\xi^{-1}\}$ and any two service range vectors $\mathsf{R},\,\mathsf{R}^{\prime}\in[0,\,\gamma]^{n}$ satisfying
@@ 2.2 Analysis for $k\geq 2$
The matching size is changed by no more than the number of supply nodes that are within distance $(\gamma/n)^{1/k}$ of a cell boundary are affected by trimming. Thus,
@@
where each constant depends only on the variables in the subscript.
@@ 2.2.1 Proof of Proposition 5
The proof follows the same structure as that of Proposition 2. We use the same notation, generalizing $\delta_{m}(r,\mathsf{R})$ to $\delta^{\varepsilon}_{m}(r,\mathsf{R})$ and $\rho_{m}^{\mathsf{R}}(r_{x},r_{y})$ to $\rho_{m}^{\mathsf{R},\varepsilon}(r_{x},r_{y})$ by replacing $G$ with $G_{\varepsilon}$ in the definitions. Throughout, trimming is performed after any augmentation: i.e. $(G\oplus x)_{\varepsilon}$ denotes the trimmed version of the augmented graph $G\oplus x$. Specifically,
[…]
and
$\rho_{m}^{\mathsf{R}}(r_{x},r_{y})\triangleq\mathbb{E}\Big[\mathds{1}\Big\{(s_{x},r_{x})\text{ and }(s_{y},r_{y})\text{ connected in }\big(G\oplus(s_{x},r_{x})\oplus(s_{y},r_{y})\big)_{\varepsilon}\Big\}\Big]\,$,
where the expectation is with respect to the randomness $s_{x},s_{y}\,\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\,\mathrm{Uniform}([0,1]^{k})$ which is independent of $G\sim\mathbb{G}(m,\mathsf{R})$. The main technical challenge is to show the success probability function $\delta_{m}^{\varepsilon}(\cdot,\mathsf{R})$ is strongly concave.
@@
For supply nodes $x=(s_{x},r_{x})$ and $y=(s_{y},r_{y})$ with $s_{x},s_{y}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Uniform}([0,1]^{k})$ independent of $G\sim\mathbb{G}(m,\mathsf{R})$ with $\mathsf{R}\in{\mathbb{R}}_{\geq 0}^{n-2}$,
$\rho_{m}^{\mathsf{R},\varepsilon}(r_{x},r_{y})\leq\beta_{k,\gamma}\,\varepsilon^{-k}/{n}\,,\text{ where }\beta_{k,\gamma}\triangleq\left(2k\right)^{k}\gamma\,.$
@@ 4. $\mathcal{M}^{\prime}\leftarrow\emptyset$ $\triangleright$ matching
5. for $t=1,\cdots,|\mathsf{D}|$ do
6.   if $d_{(t)}$ has a neighbor in $\mathsf{S}^{\prime}$ then
7.    $t^{*}\leftarrow\mathop{\arg\min}\{s_{i}+r_{i}/n:s_{i}\text{ is a neighbor of $d_{(t)}$ in $\mathsf{S}^{\prime}$}\}$ $\triangleright$ greedy matching rule
8.    $\mathsf{S}^{\prime}\leftarrow\mathsf{S}^{\prime}-s_{t^{*}}$
9.    $\mathsf{D}^{\prime}\leftarrow\mathsf{D}^{\prime}-d_{(t)}$
10.    $\mathcal{M}^{\prime}\leftarrow\mathcal{M}^{\prime}+(d_{(t)},s_{t^{*}})$
@@ 4 Simulations
In this section, we present simulations illustrating three main findings. First, we confirm that the uniformity principle holds when the service range parametrizes the volume of the service region. Second, we demonstrate that the uniformity principle no longer holds when the service range parametrizes the radius of the service region. Third, we verify the agreement between the matching rate obtained by analyzing the embedded Markov chain and the true fraction of matched nodes in a maximum matching. We then validate the upper and lower bounds (7) and (8) for the dual service range model against the simulation data.
Throughout, we generate $n=400$ supply nodes and $m=400$ demand nodes i.i.d. uniformly on $0,1]^{k}$, form the induced bipartite graph under the specified connectivity rule, and compute the maximum matching size. Curves report the average matching fraction, and shaded regions indicate $\pm 1$ standard deviation over $10^{4}$ independent trials.
To generate service range vectors that are ordered according to majorization, we use a randomized version of our dual service range model, parameterized by a scalar $\alpha\in[0,1]$:
$\mathsf{R}_{i}(\alpha)=r(\alpha)+b(\alpha)\,X_{i}(\alpha),~~~~~~X_{i}(\alpha)\stackrel{{\scriptstyle\mathrm{i.i.d}}}{{\sim}}\mathrm{Bernoulli}\big(p(\alpha)\big)\,$, (46)
i.e., each supply node has service range $r(\alpha)$ with probability $1-p(\alpha)$, and range $r(\alpha)+b(\alpha)$ with probability $p(\alpha)$. To isolate the effect of distribution shape, the parameters $r(\alpha)$, $b(\alpha)$, and $p(\alpha)$ are constrained such that the expected service range
$\bar{r}\triangleq\mathbb{E}[\mathsf{R}_{i}(\alpha)]=r(\alpha)+p(\alpha)b(\alpha)$
remains constant across all $\alpha$. The scalar $\alpha$ serves as a control for inequality: for a two-value distribution with a fixed mean, increasing $\alpha$ increases the variance, which results in a more unequal allocation, i.e. a more majorized vector.
Accordingly, we consider three families of service range vectors, each parameterized by $\alpha$: (i) the fixed $r$ family with $r(\alpha)=r_{0}$, (ii) the fixed $b$ family with $b(\alpha)=b_{0}$, and (iii) the fixed $p$ family with $p(\alpha)=p_{0}$. For each family, the varying parameters are tuned to vary monotonically with $\alpha$ such that the variance of $\mathsf{R}_{i}(\alpha)$ increases with $\alpha$. This construction ensures that
$\alpha_{1}<\alpha_{2}\implies\mathsf{R}(\alpha_{1})\preceq_{\mathrm{st}}\mathsf{R}(\alpha_{2})\,$,
i.e., a larger $\alpha$ typically produces a more majorized vector of service ranges. For example, in the fixed $r$ family with $r(\alpha)=r_{0}$, this is achieved by decreasing $p(\alpha)$ while increasing $b(\alpha)$ as $\alpha$ increases. The specific definitions for all three families are provided in Table 2.
Family Varied in $\alpha$ Implied by $\bar{r}=r+pb$
Fixed $r$: $r(\alpha)=r_{0}$ $p(\alpha)=p_{\max}-\alpha\big(p_{\max}-p_{\min}\big)$ $b(\alpha)=(\bar{r}-r_{0})/p(\alpha)$
Fixed $b$: $b(\alpha)=b_{0}$ $p(\alpha)=p_{\min}+\alpha\big(p_{\max}-p_{\min}\big)$ $r(\alpha)=\bar{r}-p(\alpha)b_{0}$
Fixed $p$: $p(\alpha)=p_{0}$ $r(\alpha)=r_{\max}-\alpha\big(r_{\max}-r_{\min}\big)$ $b(\alpha)=(\bar{r}-r(\alpha))/p_{0}$
@@
We evaluate the uniformity principle by plotting the average matching fraction (over $10^{4}$ trials) against the inequality parameter $\alpha$. We perform this evaluation for the three service range families (fixed $r$, fixed $b$, and fixed $p$) defined in Table 2. The specific numerical parameters used for dimensions $k=1,2,3$ are detailed in Table 3.
Dimension $k$ Family Mean $\bar{r}$} Parameters
11 Fixed $r$: $r(\alpha)=r_{0}$ 11 $r_{0}=0.5,\ p_{\min}=0.2,\ p_{\max}=0.8$
Fixed $b$: $b(\alpha)=b_{0}$ $b_{0}=0.6,\ p_{\min}=0.01,\ p_{\max}=0.5$
Fixed $p$: $p(\alpha)=p_{0}$ $p_{0}=0.3,\ r_{\min}=0.5,\ r_{\max}=1$
22 Fixed $r$: $r(\alpha)=r_{0}$ $0.15$ $r_{0}=0.025,\ p_{\min}=0.2,\ p_{\max}=0.8$
Fixed $b$: $b(\alpha)=b_{0}$ $b_{0}=0.3,\ p_{\min}=0.0,\ p_{\max}=0.5$
Fixed $p$: $p(\alpha)=p_{0}$ $p_{0}=0.5,\ r_{\min}=0.006,\ r_{\max}=0.12$
33 Fixed $r$: $r(\alpha)=r_{0}$ $0.1$ $r_{0}=0.02,\ p_{\min}=0.2,\ p_{\max}=0.8$
Fixed $b$: $b(\alpha)=b_{0}$ $b_{0}=0.2,\ p_{\min}=0.0,\ p_{\max}=0.5$
Fixed $p$: $p(\alpha)=p_{0}$ $p_{0}=0.5,\ r_{\min}=0.005,\ r_{\max}=0.1$
Table 3: Numerical choices used to generate $\mathsf{R}(\alpha)$ in Figure 11. For each dimension, the mean constraint $\bar{r}=r+pb$ is fixed across families, and the remaining parameters vary linearly with $\alpha$ as specified in Table 2.
Figure 11 demonstrates that for all three families and across all tested dimensions, the matching fraction decreases monotonically with $\alpha$. This observation validates the uniformity principle: a smaller $\alpha$ corresponds to a less majorized (more uniform) service range vector, which systematically yields a larger maximum matching.
@@
Figure 11: Illustration of the uniformity principle. Each plot shows three curves corresponding to the fixed $r$ with $r(\alpha)=r_{0}$, fixed $b$ with $b(\alpha)=b_{0}$, and fixed $p$ with $p(\alpha)=p_{0}$ families respectively.
Comparing volume and radius parametrizations.
As discussed in Remark 1, we investigate an alternative modeling choice where the service range parametrizes the radius of the service region rather than its volume. We focus on dimension $k=2$, where the two models are defined as follows:
  • (Volume parametrization) Given a vector $\mathsf{R}=(r_{i})_{i=1}^{n}$ of service ranges, $s_{i}$ and $d_{j}$ are connected iff
    $\|s_{i}-d_{j}\|_{2}\,\leq\,\sqrt{r_{i}/n}\,.$
  • (Radius parametrization) Given a vector $\mathsf{R}=(r_{i})_{i=1}^{n}$ of service ranges, $s_{i}$ and $d_{j}$ are connected iff
    $\|s_{i}-d_{j}\|_{2}\leq r_{i}/\sqrt{n}\,.$
We evaluate both models using the same three families defined in Table 2 (fixed $r$, fixed $b$, and fixed $p$). The specific numerical parameters for these simulations are detailed in Table 4. The matching fractions for both models are plotted in Figure 12. Consistent with Theorem 1, the volume parametrization obeys the uniformity principle, evidenced by the matching fraction decreasing monotonically with $\alpha$ across all panels. In contrast, the radius model violates this principle; the matching fraction is not monotonically decreasing. Specifically, there exist instances where $\alpha_{1}<\alpha_{2}$ (implying $\mathsf{R}(\alpha_{1})\preceq_{\mathrm{st}}\mathsf{R}(\alpha_{2})$) yet the matching fraction is strictly larger at $\alpha_{2}$.
(a) Fixed $r$ with $r(\alpha)=r_{0}$
(b) Fixed $b$ with $b(\alpha)=b_{0}$
(c) Fixed $p$ with $p(\alpha)=p_{0}$
Figure 12: Illustration of the uniformity principle holding when service range parametrizes volume, but not holding when it parametrizes radius, for the three families in dimension $k=2$ and parameters in Table 4.
(a) Varying $r$
(b) Varying $b$
(c) Varying $p$
@@
Family Radius interpretation Volume interpretation
Fixed $r$ $\bar{r}=0.3,\,r_{0}=0,\,p\in[0.2,0.8]$ $\bar{r}=0.127,\,r_{0}=0,\,p\in[0.2,0.8]$
Fixed $b$ $\bar{r}=0.4,\,b_{0}=0.8,\,p\in[0,0.5]$ $\bar{r}=0.15,\,b_{0}=0.09,\,p\in[0,0.5]$
Fixed $p$ $\bar{r}=0.4,\,p_{0}=0.5,\,r\in[0.02,0.36]$ $\bar{r}=0.21,\,p_{0}=0.5,\,r\in[0.011,\,0.191]$
@@
Figure 13 checks the accuracy of the closed-form expression (2), which is derived from the stationary distribution of the Markov chain $\boldsymbol{\psi}(t)$. For each parameter setting, we plot (i) the matching fraction obtained by the embedded Markov chain, and (ii) the exact matching fraction obtained by solving the underlying bipartite matching instance via the Hopcroft–Karp algorithm. We plot the two quantities under three separate parameter sweeps, varying $r$, $b$, and $p$ (holding the remaining parameters fixed). In all panels, the theoretical and empirical curves agree, validating the formula (2).
Finally, Figure 14 verifies the theoretical bounds derived in Remark 4 for the dual-service range model. We plot these bounds as functions of $r$, $b$, and $p$ (holding other parameters fixed) and observe that they effectively bracket the true matching fraction. The upper bound (7) corresponds to the performance of the uniform allocation (via the uniformity principle), while the lower bounds (8) and (9) represent the performance guaranteed by decomposing the graph into disjoint subgraphs.